Perform transformation of stress for a point in a body, for which σxx = 90 MPa; σyy = -10 MPa; τxy = 40 MPa.

a. Find tensor components for a point P on a cutting plane Q with normal vector n = 1/sqrt(2)i - 1/sqrt(2)j

b. Find the state of stress for an element rotated by θ = π/6 clockwise (negative direction) from x-axis.
Answer #1

Sfaass tenSO 40- lo Ca Ceting ptane VE 40-40 ·ơMpa.= 40-0+4 Scn 50 50 80 cos(-Go+ 40十25-2048 Tyl = 15+2013 MPa 65-03 20+25 SissG Tensor a 20 +25 54 203